OpenStudy (anonymous):

Find the area under the curve from y= cosx/ pi +sinx from 0 to infiniti

OpenStudy (anonymous):

What math you take?

OpenStudy (anonymous):

Calculus :/

OpenStudy (anonymous):

Your a Junior?

OpenStudy (anonymous):

Integrate form 0 to infinity

OpenStudy (anonymous):

\[\frac{\text{Cos}[x]}{\pi +\text{Sin}[x]}\] did I write you function correctly?

OpenStudy (anonymous):

senior. imranmeah can you help me out a little more buddy? not exactly im stuck between 0, 2.76, .693, and .079 :/ please help! kay thanks :)

OpenStudy (anonymous):

did imranmeah91 wrote your function correctly?

OpenStudy (anonymous):

Yes, I am wondering if I copies the equation correctly above?

OpenStudy (anonymous):

yeahhh :)

OpenStudy (anonymous):

its right so far...

OpenStudy (anonymous):

\[\int_0^{\infty } \frac{\cos (x)}{\sin (x)+\pi } \, dx\] Simple U-Sub from here u=Sin(x)+pi du= cos(x) dx Plug back in

OpenStudy (anonymous):

\[\int_0^{\infty } \frac{1}{u} \, du\]

OpenStudy (anonymous):

speak for yourself. :x this is like a whole different language to me :/

OpenStudy (anonymous):

How much have you learned in regard to integration?

OpenStudy (anonymous):

Welp to be honest, not much. This is complete jargon. That's why I need your help! Kay thanks :)

OpenStudy (anonymous):

ln(sin(x)+pi)|[0,inf] ln(sin(inf)+pi)-ln(pi) sin(inf) isnt defined its always going back to +1---->-1 to reason it out i would say since ln is only positive it would look like ln(1+pi)-ln(pi) .2763

OpenStudy (anonymous):

or twice that sin from o to pi is 2 and if it never goes negative then you would double that, but this question has something wrong because sin(inf)is not defines in calc 1 the lim as x--->inf of sin(x) is not defined because it doesnt hit anything